I am really confused at the AC concept. Can someone please help? Thanks!
Use the ac test to determine if x^2 +2x-15 can be factored. Find the values of m and n if it can be factored.
Use the ac test to determine if x^2-3x+7 can be factored. Find the values of m and n if it can be factored.

Answer by stanbon(75887)

Use the ac test to determine if x^2 +2x-15 can be factored. Find the values of m and n if it can be factored.
--------------------
Think of two numbers whose product is A*C=-15 and whose sum = B= 2
The numbers are 5 and -3
Replace 2x with 5x-3x in the original problem:
x^2+5x-3x-15
Factor the 1st two and the last two terms separately:
x(x+5)-3(x+5)
Factor again to get the final answer:
(x+5)(x-3)
========================
Use the ac test to determine if x^2-3x+7 can be factored. Find the values of m and n if it can be factored.
Think of two numbers whose product is AC = 1*7 = 7 and whose sum = B= -3
There are no such numbers so this quadratic cannot factored by the AC Method.
------------
Cheers,
Stan H.